# Stronger, Fewer, & Superior: Harnessing Vision Foundation Models for Domain Generalized Semantic Segmentation

### v4

@@ Abstract
In this paper, we first assess and harness various Vision Foundation Models (VFMs) in the context of Domain Generalized Semantic Segmentation (DGSS). Driven by the motivation that Leveraging Stronger pre-trained models and Fewer trainable parameters for Superior generalizability, we introduce a robust fine-tuning approach, namely “Rein”, to parameter-efficiently harness VFMs for DGSS. Built upon a set of trainable tokens, each linked to distinct instances, Rein precisely refines and forwards the feature maps from each layer to the next layer within the backbone. This process produces diverse refinements for different categories within a single image. With fewer trainable parameters, Rein efficiently fine-tunes VFMs for DGSS tasks, surprisingly surpassing full parameter fine-tuning. Extensive experiments across various settings demonstrate that Rein significantly outperforms state-of-the-art methods. Remarkably, with just an extra 1% of trainable parameters within the frozen backbone, Rein achieves a mIoU of 68.1% on the Cityscapes, without accessing any real urban-scene datasets. Code is available at https://github.com/w1oves/Rein.git.
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Table 1: Performance benchmarking of multiple VFMs and previous DGSS methods under the GTAV normal-$\rightarrow$ Cityscapes (Citys) + BDD100K (BDD) + Mapillary (Map) generalization setting. Models are fine-tuned on GTAV and tested on Cityscapes, BDD100K and Mapillary. The best results are highlighted. Without specialized design, frozen VFMs demonstrate significantly stronger performance.
@@ 1 Introduction
Fewer: Although VFMs have exhibited impressive generalization capabilities, fine-tuning them for DGSS tasks poses a challenge. The datasets [53, 10] commonly used in DGSS tasks are significantly smaller in scale compared to ImageNet [11], and fine-tuning VFMs with their huge number of trainable parameters on these datasets result in limited generalizability [27]. To address this issue, instead of the difficult task of large datasets collection, we resort to fine-tuning VFMs with fewer trainable parameters. However, most existing parameter-efficient fine-tuning strategies, which fine-tune a large-scale model with fewer trainable parameters, are primarily designed for adapting large language models [21, 66, 20, 36, 38, 40, 71] or classification networks [5, 23]. These methods often lack precision in refining features for distinct instances within a single image, thereby limiting their effectiveness in DGSS contexts.
Superior: In this work, we introduce a robust and efficient fine-tuning approach, namely “Rein”. Tailored for DGSS tasks, Rein employs fewer trainable parameters to harness stronger VFMs for achieving superior generalization. At its core, Rein comprises a set of randomly initialized tokens, each directly linked to different instances. These tokens, through a dot-product operation with VFMs features, generate an attention-like similarity map. This map enables Rein to perform precise refinement tailored to each instance within an image, significantly boosting VFMs in the context of DGSS. Moreover, to reduce the number of trainable parameters, we employ shared weights across MLPs in different layers and design our learnable tokens by multiplying two low-rank matrices. Extensive experiments on various DGSS settings demonstrate that the proposed Rein outperforms existing DGSS methods by a large margin with fewer trainable parameters. In a nutshell, the main contributions of this paper are as follows:
@@
  • We present a robust fine-tuning method, namely “Rein”, to parameter-efficiently harness VFMs. At its core, Rein consists of a set of learnable tokens, each directly linked to distinct instances. With deliberate design, this linkage enables Rein to refine feature maps at an instance-level within each backbone layer. As a result, Rein reinforces the ability of VFMs in DGSS tasks, achieving this with Fewer trainable parameters while preserving the pre-trained knowledge.
@@
Figure 2: An overview of proposed Rein. Rein primarily consists of a collection of low-rank learnable tokens, denoted as $T=\{T_{1},T_{2},\ldots,T_{N}\}$. These tokens establish direct connections to distinct instances, facilitating instance-level feature refinement. This mechanism results in the generation of an enhancement feature map $f^{\prime}_{i}=f_{i}+Rein(f_{i})$ for each layer within backbone. All MLPs share same parameters to reduce the number of parameters. The notation $max\leavevmode\nobreak\ \&\leavevmode\nobreak\ avg\leavevmode\nobreak\ \&\leavevmode\nobreak\ last$ refers to the equation Eq. (8) and Eq. (10).
2 Related Works
Domain Generalized Semantic Segmentation. Domain Generalized Semantic Segmentation (DGSS) focuses on enhancing model generalizability. This field typically involves training models on a set of source domain data to enhance their performance on distinct and unseen target domain datasets. Various approaches [12, 22, 24, 25, 50] have been proposed to address this issue within DGSS, with representative methods including splitting the learned features into domain-invariant and domain-specific components [63, 59], or employing meta-learning to train more robust models [29]. A standard scenario in DGSS is generalizing from one urban-scene dataset to another, for instance, from the synthetic GTAV [53] dataset to the real-world Cityscapes [10]. In this classic setting, certain techniques [8, 47, 48] have achieved notable performance through learning feature normalization/whitening schemes, while others [35] have further improved segmentation results through feature-level style transfer and the introduction of additional data. Additionally, strong data augmentation [69, 49, 4] often simply and effectively enhances model robustness. However, most of previous DGSS methods generally utilize outdated backbones like ResNet [18], VGGNet [58], MobileNetV2 [55], and ShuffleNetV2 [43], thereby leaving the efficacy of stronger Vision Foundation Models (VFMs) in DGSS relatively unexplored.
Vision Foundation Models. The concept of a Foundation Model, initially introduced by Bommasani et al. [1] in the field of Natural Language Processing (NLP), defined as “the base models trained on large-scale data in a self-supervised or semi-supervised manner that can be adapted for several other downstream tasks”. While models like the ViT [13] and Swin Transformer [41] have demonstrated excellent performance, the quest for a Vision Foundation Model (VFM) similar to their NLP counterparts is ongoing. This pursuit has yielded significant advancements with the advent of models such as CLIP [51], which learn high-quality visual representation by exploring contrastive learning with large-scale image text pairs; MAE [19], utilizing a masked image modeling framework for learning latent image representations; SAM [33], which develops a promptable model and pre-train it on a broad dataset for segmentation task; EVA02 [15, 14], which integrates Masked Image Modeling pre-training with CLIP’s vision features as the pretext task’s target; and DINOv2 [46], which is pretrained on extensive, carefully curated datasets without explicit supervision. These VFMs have shown remarkable performance in a variety of downstream applications, demonstrating their impressive generalization capabilities. Yet, a dedicated investigation into their performance in the specific context of DGSS tasks remains unexplored.
Parameter-Efficient Fine-tuning. In the NLP domain, parameter-efficient fine-tuning (PEFT) has achieved notable success by freezing most parameters of the foundation model and fine-tuning a select few. Various strategies have been introduced, such as BitFit [66], which tweaks only the bias-terms of the model, or just a subset of these terms; Prompt-tuning [36], which learns soft prompts to condition frozen language models to perform specific downstream tasks; Adapter-tuning [20], which incorporates extra lightweight modules within each Transformer layer; and notably, LoRA [21], which injects trainable rank decomposition matrices into each layer of transformer architecture, yielding significant influence. PEFT methods are gaining traction in computer vision as well, exemplified by Visual Prompt Tuning [23], which prepends prompts into the input sequence of Transformer layers for fine-tuning, and AdaptFormer [5], which replaces MLP block in the transformer encoder with an AdaptMLP comprising two sub-branches. However, these methodologies are primarily tuned for classification tasks, where each image contains only one target to identify. Our endeavor is tailored for segmentation tasks, refining feature maps at the object-level for each instance in the image, thereby achieving superior performance.
Backbone Fine-tune Method Trainable $*{}^{*}$ mIoU
@@ 2 Related Works
Table 2: Performance Comparison with the proposed Rein across Multiple VFMs as Backbones under the GTAV normal-$\rightarrow$ Cityscapes (Citys) + BDD100K (BDD) + Mapillary (Map) generalization setting. Models are fine-tuned on GTAV and tested on Cityscapes, BDD100K and Mapillary. The best results are highlighted. *** denotes trainable parameters in backbones.
Backbone Fine-tune Method Trainable $*{}^{*}$ mIoU
@@
Table 3: Performance Comparison of the proposed Rein against other DGSS and PEFT methods under the GTAV normal-$\rightarrow$ Cityscapes (Citys) + BDD100K (BDD) + Mapillary (Map) generalization setting. Models are fine-tuned on GTAV and tested on Cityscapes, BDD100K and Mapillary. The best results are highlighted. *** denotes trainable parameters in backbones.
Target ACDC (test) All
Night Snow Fog Rain
HGFormer 52.7 68.6 69.9 72.0 67.2
Ours 70.6 79.5 76.4 78.2 77.6
Table 4: Performance Comparison on the ACDC test set.
Target Blur Noise Digital Weather Avg.
HGFormer 64.1 67.2 61.5 63.6 27.2 35.7 32.9 63.1 79.9 72.9 78.0 53.6 55.4 75.8 75.5 43.2 59.4
Ours 68.5 71.7 69.7 68.7  6.2 23.0 13.1 63.7 81.5 78.9 80.6 68.8 63.8 73.6 79.5 47.9 60.0
Table 5: Performance Comparison on the Cityscapes-C validation set.
Backbone
ResNet
ResNet
ConvNeXt
ViT
ViT
Full 38.9 46.1 52.2 51.8 56.7
Ours 46.6 46.3 55.5 55.7 59.1
Table 6: Results for ConvNets and smaller backbones.
@@ 3.1 Preliminary
Driven by the motivation that Leveraging Stronger pre-trained models and Fewer trainable parameters for Superior generalizability, we choose to fine-tune VFMs with a reduced parameter set. A straightforward thought might involve a smaller decode head; however, this method merely acts as a passive receiver of feature maps from the backbone, lacking the flexibility to effectively adapt a frozen backbone for generating task-specific or scene-specific features. In contrast, we propose to embed a mechanism, named “Rein”, between the layers within the backbone. Rein actively refines and forwards the feature maps from each layer to the subsequent one. This approach allows us to more effectively utilize the powerful capabilities of VFMs, much like using rein to control a horse.
@@
where $x_{i}$ and $y_{i}$ denote the input image and its corresponding ground truth, respectively, and $N_{d}$ signifies the total number of datasets. $\mathcal{F}_{\Phi_{M},\theta_{r}}$ represents the forward process of VFM after applying the Rein strategy.
@@ 3.2 Core of Rein
where $f^{\prime}_{i}=f_{i}+\Delta f_{i}$ symbolizes the refined feature map, $x$ is the input image, $Embed$ denotes the patch embedding layer in VFMs, $n$ represents the number of patches, and $c$ is the dimensionality of $f_{1},f_{2},\ldots,f_{N}$. Note that the layers $L_{1},L_{2},\ldots,L_{N}$ are kept frozen, and our focus is on training an efficient module, Rein, to generate $\Delta f_{i}$ as follows:
@@
where $W_{T_{i}}$ and $b_{T_{i}}$ denote the weights and biases of a MLP, respectively. This MLP enables the transformation of $T_{i}$ across different feature spaces during the computation of $S_{i}$ and $\Delta\bar{f_{i}}$. Optionally, Rein can pre-calculate $T_{i}\times W_{T_{i}}+b_{T_{i}}$ to reduce inference time. Notably, $S_{i}(:,2:m)$ selects columns 2 to m of $S_{i}$, and $T_{i}(2:m)$ denotes the selection of rows 2 to m of $T_{i}$. This selection is particularly useful in handling challenging samples that might not have corresponding tokens in $T_{i}$. In these cases, the total similarity in the respective row of $S_{i}$ remains $1$, potentially leading to erroneous modifications. To counter this, Rein excludes the first token of $T_{i}$ and the first column of $S_{i}$, enabling the sum of each row in $S_{i}$ to range between 00 and $1$, thereby reducing the risk of inappropriate alterations.
@@ 4.1 Settings
Visual Foundation Models. To thoroughly assess the influence of Visual Foundation Models (VFMs) within the context of DGSS, we analyze five distinct VFMs, each with different training strategies and datasets. Our selection includes CLIP [51], a language-image pre-training model; MAE [19], known for its masked pre-training approach; SAM [33], which leverages a large-scale segmentation dataset; EVA02 [15, 14] combines CLIP with masked image modeling; and DINOv2 [46], based on self-supervised pre-training with curated dataset. For balancing precision and efficiency, we mainly employ the ViT-Large architecture for these VFMs, except SAM, which utilizes a ViT-Huge image encoder, as described in its original paper [33]. We establish two fundamental baselines for VFMs: “Full”, where we fine-tune the entire network, and “Freeze”, in which all backbone parameters are fixed, with training solely on the segmentation head. More details about VFMs are available in the supplementary material.
Methods Publication mIoU
Citys BDD Map Avg.
RobustNet [8] CVPR 21 37.7 34.1 38.5 36.8
PintheMem [29] CVPR 22 44.5 38.1 42.7 41.8
SAN-SAW [50] CVPR 22 42.1 37.7 42.9 40.9
WildNet [35] CVPR 22 43.7 39.9 43.3 42.3
DIGA [61] CVPR 23 46.4 33.9 43.5 41.3
SPC [22] CVPR 23 46.4 43.2 48.2 45.9
EVA02 - Frozen [15, 14] arXiV 23 55.8 55.1 59.1 56.7
EVA02 + Rein - 63.5 60.7 63.9 62.7
DINOv2 - Frozen [46] arXiV 23 64.8 60.2 65.2 63.4
DINOv2 + Rein - 68.1 60.5 67.1 65.2
Table 7: Performance Comparison of the proposed Rein against other DGSS methods under GTAV + Synthia normal-$\rightarrow$ Cityscapes (Citys) + BDD100K (BDD) +Mapillary (Map) generalization. Models are fine-tuned on GTAV and Synthia, tested on Cityscapes, BDD100K and Mapillary. The best results are highlighted.
Methods Backbone Trainable $*{}^{*}$ mIoU
BDD Map Avg.
IBN [47] ResNet50 [18]   23.58M 48.6 57.0 52.8
DRPC [65] ResNet50 [18]   23.58M 49.9 56.3 53.1
GTR [49] ResNet50 [18]   23.58M 50.8 57.2 54.0
SAN-SAW [50] ResNet50 [18]   23.58M 53.0 59.8 56.4
WildNet [35] ResNet101 [18]   42.62M 50.9 58.8 54.9
HGFormer [12] Swin-L [41] 196.03M 61.5 72.1 66.8
Freeze EVA02-L [14]    0.00M 57.8 63.8 60.8
Rein (Ours) EVA02-L [14]    2.99M 64.1 69.5 66.8
Freeze DINOv2-L [46]    0.00M 63.4 69.7 66.7
Rein (Ours) DINOv2-L [46]    2.99M 65.0 72.3 68.7
Table 8: Performance Comparison of the Rein against other DGSS methods under Cityscapes normal-$\rightarrow$ BDD100K (BDD) +Mapillary (Map) generalization. Models are fine-tuned on Cityscapes and tested on BDD and Map. The best results are highlighted.
@@
Table 9: Ablation Study about Rein under Cityscapes normal-$\rightarrow$ BDD100K generalization in terms of mIoU. Components are sequentially incorporated. To better illustrate the gains contributed by each component, we employ varying shades of yellow to demonstrate the relative performance of the Freeze and Rein methods. The best results across all methods are highlighted.
Figure 3: Qualitative Comparison under GTAV normal-$\rightarrow$ Cityscapes (Citys) + BDD100K (BDD) + Mapillary (Map) generalization setting.
Datasets. We evaluate VFMs and proposed methods on both real-world datasets (Cityscapes [10], BDD100K [64], Mapillary [45]) and synthetic datasets (GTAV [53], Synthia [54]). In detail, Cityscapes (denoted as Citys) is an autonomous driving dataset that contains 2975 training images and 500 validation images, each with the resolution of $2048\times 1024$. BDD100K (shortened to BDD) and Mapillary (denoted by Map) offer 1,000 ($1280\times 720$) and 2,000 ($1902\times 1080$) validation images, respectively. GTAV, a synthetic dataset, presents 24,966 labeled images obtained from the game. Synthia, another synthetic dataset, provides 25,000 images created by photo-realistic rendering.

### v5

@@ Abstract
In this paper, we first assess and harness various Vision Foundation Models (VFMs) in the context of Domain Generalized Semantic Segmentation (DGSS). Driven by the motivation that Leveraging Stronger pre-trained models and Fewer trainable parameters for Superior generalizability, we introduce a robust fine-tuning approach, namely “Rein", to parameter-efficiently harness VFMs for DGSS. Built upon a set of trainable tokens, each linked to distinct instances, Rein precisely refines and forwards the feature maps from each layer to the next layer within the backbone. This process produces diverse refinements for different categories within a single image. With fewer trainable parameters, Rein efficiently fine-tunes VFMs for DGSS tasks, surprisingly surpassing full parameter fine-tuning. Extensive experiments across various settings demonstrate that Rein significantly outperforms state-of-the-art methods. Remarkably, with just an extra 1% of trainable parameters within the frozen backbone, Rein achieves a mIoU of 78.4% on the Cityscapes, without accessing any real urban-scene datasets. Code is available at https://github.com/w1oves/Rein.git.
[Uncaptioned image] [Uncaptioned image]
@@
Table 1: Performance benchmarking of multiple VFMs and previous DGSS methods under the GTAV $\rightarrow$ Cityscapes (Citys) + BDD100K (BDD) + Mapillary (Map) generalization setting. Without specialized design, frozen VFMs demonstrate stronger performance.
@@ 1 Introduction
Fewer: Although VFMs have exhibited impressive generalization capabilities, fine-tuning them for DGSS tasks poses a challenge. The datasets [65, 12] commonly used in DGSS tasks are significantly smaller in scale compared to ImageNet [13], and fine-tuning VFMs with their huge number of trainable parameters on these datasets result in limited generalizability [36]. To address this issue, instead of the difficult task of large datasets collection, we resort to fine-tuning VFMs with fewer trainable parameters. However, most existing parameter-efficient fine-tuning strategies, which fine-tune a large-scale model with fewer trainable parameters, are primarily designed for adapting large language models [29, 80, 28, 46, 49, 51, 85] or classification networks [7, 31]. These methods are not developed for refining features for distinct instances within a single image, thereby limiting their effectiveness in DGSS contexts.
Superior: In this work, we introduce a robust and efficient fine-tuning approach, namely “Rein". Tailored for DGSS tasks, Rein employs fewer trainable parameters to harness stronger VFMs for achieving superior generalization. At its core, Rein comprises a set of randomly initialized tokens, each directly linked to different instances. These tokens, through a dot-product operation with VFMs features, generate an attention-like similarity map. This map enables Rein to perform precise refinement tailored to each instance within an image, significantly boosting VFMs in the context of DGSS. Moreover, to reduce the number of trainable parameters, we employ shared weights across MLPs in different layers and design our learnable tokens by multiplying two low-rank matrices. Extensive experiments on various DGSS settings demonstrate that the proposed Rein outperforms existing DGSS methods by a large margin with fewer trainable parameters. In a nutshell, the main contributions of this paper are as follows:
@@
  • We present a robust fine-tuning method, namely “Rein", to parameter-efficiently harness VFMs. At its core, Rein consists of a set of learnable tokens, each directly linked to instances. With deliberate design, this linkage enables Rein to refine features at an instance-level within each layer. As a result, Rein reinforces the ability of VFMs in DGSS tasks, achieving this with Fewer trainable parameters while preserving the pre-trained knowledge.
@@
Figure 2: An overview of proposed Rein. Rein primarily consists of a collection of low-rank learnable tokens, denoted as $T=\{T_{1},T_{2},\ldots,T_{N}\}$. These tokens establish direct connections to distinct instances, facilitating instance-level feature refinement. This mechanism results in the generation of an enhancement feature map $f^{\prime}_{i}=f_{i}+Rein(f_{i})$ for each layer within backbone. All parameters of MLPs are layer-shared to reduce the number of parameters. $M_{f}$, $M_{Q}$, and $M_{S}$ are features module, queries module, and segmentation module, respectively. The notation $max\leavevmode\nobreak\ \&\leavevmode\nobreak\ avg\leavevmode\nobreak\ \&\leavevmode\nobreak\ last$ refers to the equation Eq. (8) and Eq. (10).
Backbone Fine-tune Method Trainable Params mIoU
@@
Table 2: Performance Comparison with the proposed Rein across Multiple VFMs as Backbones under the GTAV $\rightarrow$ Cityscapes (Citys) + BDD100K (BDD) + Mapillary (Map) generalization setting. Mark * denotes trainable parameters in backbones.
2 Related Works
Domain Generalized Semantic Segmentation. Domain Generalized Semantic Segmentation (DGSS) focuses on enhancing model generalizability. This field involves training models on a set of source domain to enhance their performance on distinct and unseen target domain. Various approaches [14, 30, 32, 33, 62, 34, 6, 75, 48, 22, 23] have been proposed to address this issue, with methods including splitting the learned features into domain-invariant and domain-specific components [77, 72], or employing meta-learning to train more robust models [38]. A standard scenario in DGSS is generalizing from one urban-scene dataset to another, for instance, from the synthetic GTAV [65] dataset to the real-world Cityscapes [12]. In this classic setting, certain techniques [10, 59, 60] have achieved notable performance through learning feature normalization/whitening schemes, while others [45] have improved segmentation results through feature-level style transfer and the introduction of additional data. Additionally, strong data augmentation [83, 61, 5, 17] often simply and effectively enhances model robustness. However, most of previous DGSS methods generally utilize outdated backbones like ResNet [26], VGGNet [71], MobileNetV2 [68], and ShuffleNetV2 [54], thereby leaving the efficacy of stronger Vision Foundation Models (VFMs) in DGSS relatively unexplored.
Vision Foundation Models. The concept of a Foundation Model, initially introduced by Bommasani et al. [2] in the field of Natural Language Processing (NLP), defined as “the base models trained on large-scale data in a self-supervised or semi-supervised manner that can be adapted for several other downstream tasks". While models like the ViT [15] and Swin Transformer [52] have demonstrated excellent performance, the quest for a Vision Foundation Model (VFM) similar to their NLP counterparts is ongoing. This pursuit has yielded significant advancements with the advent of models such as CLIP [63], which learn high-quality visual representation by exploring contrastive learning with large-scale image text pairs; MAE [27], utilizing a masked image modeling for learning latent image representations; SAM [42], which develops a promptable model and pre-train it on a broad dataset for segmentation task; EVA02 [19, 18], which integrates Masked Image Modeling pre-training with CLIP’s vision features; and DINOv2 [58], which is pretrained on extensive, curated datasets without explicit supervision. These VFMs have shown remarkable performance in downstream applications. Yet, a dedicated investigation into their performance in the specific context of DGSS tasks remains unexplored.
Parameter-Efficient Fine-tuning. In the realm of NLP, parameter-efficient fine-tuning (PEFT) has achieved notable success by freezing most parameters of VFMs and fine-tuning a select few. Various approaches have been developed, such as BitFit [80], which adjusts only the model’s bias terms; Prompt-tuning [46], introducing soft prompts to adapt frozen language models; Adapter-tuning [28], adding lightweight modules to each transformer layer; and notably, LoRA [29], which injects trainable rank decomposition matrices into transformer layers, yielding significant influence.
The application of PEFT methods is also expanding into the field of computer vision [16, 43], with notable examples such as Visual Prompt Tuning (VPT) [31], which prepends prompts into the input sequence of transformer layers; AdaptFormer [7], replacing the MLP block in the transformer encoder with an AdaptMLP; LP-FT [43] find that fine-tuning can achieve worse accuracy than linear probing out-of-distribution; and Prompt-ICM [20], applying large-scale pre-trained models to the task of image coding for machines. Contrasting with these methods, we aim to refine feature maps for each instance within an image, thereby achieving superior performance in the realm of DGSS.
Target ACDC[67] (test) Cityscapes-C[56] (level-5)
Night Snow Fog Rain All Blur Noise Digital Weather Avg.
HGFormer 52.7 68.6 69.9 72.0 67.2 64.1 67.2 61.5 63.6 27.2 35.7 32.9 63.1 79.9 72.9 78.0 53.6 55.4 75.8 75.5 43.2 59.4
Ours 70.6 79.5 76.4 78.2 77.6 68.5 71.7 69.7 68.7  6.2 23.0 13.1 63.7 81.5 78.9 80.6 68.8 63.8 73.6 79.5 47.9 60.0
Table 3: Results on Cityscapes $\rightarrow$ ACDC (test) and Cityscapes-C (level-5) datasets, utilizing a batch size of 8.
Backbone Fine-tune Method Trainable Params mIoU
@@ 2 Related Works
Table 4: Performance Comparison of the proposed Rein against other DGSS and PEFT methods under the GTAV $\rightarrow$ Cityscapes (Citys) + BDD100K (BDD) + Mapillary (Map) generalization setting. Mark * denotes trainable parameters in backbones.
Source Domain Cityscapes mIoU
GTAV 66.4
+Synthia 68.1
+UrbanSyn 78.4
+1/16 of Cityscapes Training set 82.5
Table 5: Synthetic data + 1/16 of Citys. $\rightarrow$ Citys. val set.
@@ 3.1 Preliminary
Driven by the motivation that Leveraging Stronger pre-trained models and Fewer trainable parameters for Superior generalizability, we choose to fine-tune VFMs with a reduced parameter set. A straightforward thought might involve a smaller decode head; however, this method merely acts as a passive receiver of feature maps from the backbone, lacking the flexibility to effectively adapt a frozen backbone for generating task-specific or scene-specific features. In contrast, we propose to embed a mechanism, named “Rein", between the layers within the backbone. Rein actively refines and forwards the feature maps from each layer to the subsequent one. This approach allows us to more effectively utilize the powerful capabilities of VFMs, much like using rein to control a horse.
@@
where $x_{i}$ and $y_{i}$ denote the input image and its corresponding ground truth, respectively, and $N_{d}$ signifies the total number of samples. $\mathcal{F}_{\Phi_{M},\theta_{R}}$ represents the forward process of VFM after applying the Rein strategy.
@@ 3.2 Core of Rein
where $f^{\prime}_{i}=f_{i}+\Delta f_{i}$ symbolizes the refined feature map, $x$ is the input image, $Embed$ denotes the patch embedding layer in VFMs, $n$ represents the number of patches, $N$ denotes the number of layers, and $c$ is the dimensionality of $f_{1},f_{2},\ldots,f_{N}$. Note that the layers $L_{1},L_{2},\ldots,L_{N}$ are kept frozen, and our focus is on training an efficient module, Rein, to generate $\Delta f_{i}$ as follows:
@@
where $W_{T_{i}}$ and $b_{T_{i}}$ denote the weights and biases of a MLP, respectively. This MLP enables the transformation of $T_{i}$ across different feature spaces during the computation of $S_{i}$ and $\Delta\bar{f_{i}}$. Optionally, Rein can pre-calculate $T_{i}\times W_{T_{i}}+b_{T_{i}}$ to reduce inference time. The sum of $S_{i}$ equals one due to the softmax function; however, this can induce unneeded changes when all features are precise. To avoid this, $S_{i}(:,2:m)$ is designed to choose columns 2 to m of $S_{i}$, and $T_{i}(2:m)$ denotes the selection of rows 2 to m of $T_{i}$. This strategic selection allows models to sidestep unnecessary adjustments by assigning a high value to the first token and subsequently discarding it. This approach allows the sum of each row in $S_{i}$ to vary from 00 to $1$, thus reducing the risk of inappropriate changes.
@@ 4.1 Settings
Visual Foundation Models. To thoroughly assess the influence of Visual Foundation Models (VFMs) within the context of DGSS, we analyze five distinct VFMs, each with different training strategies and datasets. Our selection includes CLIP [63], a language-image pre-training model; MAE [27], known for its masked pre-training approach; SAM [42], which leverages a large-scale segmentation dataset; EVA02 [19, 18] combines CLIP with masked image modeling; and DINOv2 [58], based on self-supervised pre-training with curated dataset. For balancing precision and efficiency, we mainly employ the ViT-Large architecture for these VFMs, except SAM, which utilizes a ViT-Huge image encoder, as described in its original paper [42]. We establish two fundamental baselines for VFMs: “Full", where we fine-tune the entire network, and “Freeze", in which all backbone parameters are fixed, with training solely on the segmentation head. More details about VFMs and PEFT methods are available in the supplementary material.
@@
Table 6: Ablation Study about Rein under Cityscapes $\rightarrow$ BDD100K generalization in terms of mIoU. Components are sequentially incorporated. To better illustrate the gains contributed by each component, we employ varying shades of yellow to demonstrate the relative performance of the Freeze and Rein methods. The best results across all methods are highlighted.
Figure 3: Qualitative Comparison under GTAV $\rightarrow$ Cityscapes (Citys) + BDD100K (BDD) + Mapillary (Map) generalization setting.
Datasets. We evaluate VFMs and proposed methods on both real-world datasets (Cityscapes [12], BDD100K [78], Mapillary [57]) and synthetic datasets (GTAV [65], Synthia [66], UrbanSyn [24]). In detail, Cityscapes (denoted as Citys) is an autonomous driving dataset that contains 2975 training images and 500 validation images, each with the resolution of $2048\times 1024$. BDD100K (shortened to BDD) and Mapillary (denoted by Map) offer 1,000 ($1280\times 720$) and 2,000 ($1902\times 1080$) validation images, respectively. GTAV, a synthetic dataset, presents 24,966 labeled images obtained from the game. Synthia, a synthetic dataset, provides 25,000 images created by photo-realistic rendering. UrbanSyn, a synthetic dataset consists of 7,539 images.